OpenStudy (anonymous):

how do i get the derivative of these two: *2e^θ *e^-θ/2

OpenStudy (yttrium):

Take note that \[\frac{ d }{ dx }(e^x) = e^x\] Therefore, \[\frac{ d }{ d \theta } (2e^\theta) = 2e^\theta\]

OpenStudy (yttrium):

And \[\frac{ d }{ d \theta } (e^-\frac{ \theta }{ 2 })= \frac{ -1 }{ 2 } (e^\frac{ -\theta }{ 2 }) \]

OpenStudy (anonymous):

Thanks!!